# Multimodal Fusion Balancing Through Game-Theoretic Regularization

### v2

@@ 4.2 Real-World Datasets
    Our method, MCR, stands out as the only approach that consistently achieves better across all datasets and backbone models. Notably, MCR outperforms consistently the Ensemble baseline across all models and datasets, demonstrating its effectiveness in balancing modality contributions under several different configurations.
@@
4.3 Ablation Study
Strategy: We perform an ablation study to test our hypothesis that framing multimodal competition regularization as a game benefits the model by avoiding destructive loss interactions in each backbone encoder. Table 2 compares the three strategies: Collaborative, Independent, and Greedy. The results show that, across the models allowing backbone encoders to maximize their own $\operatorname{CMI}$ term and concurrently minimizing the others (Greedy strategy) consistently yields the best performance. This result demonstrates that framing multimodal models as competing modalities using game-theoretic principles in the loss terms can be beneficial in balancing these loss terms.
Table 2: Ablation Study: Game Strategies – Comparison of model accuracy for different game strategies: Collaborative, Independent, and Greedy, using both ResNet and Conformer backbones.
Collaborative Independent Greedy
Model Setting $\min\limits_{\theta}\mathcal{L}_{\operatorname{MIPD}}$ $\min\limits_{\theta_{i}}\mathcal{L}_{\operatorname{MIPD_{X_{i}}}}$ $\min\limits_{\theta_{i}}\max\limits_{\theta_{\neg i}}\mathcal{L}_{\operatorname{MIPD_{X_{i}}}}$
ResNet + MCR 73.4$\pm$3.0 76.0$\pm$2.0 76.2$\pm$1.7
Conformer + MCR 82.9$\pm$0.7 82.9$\pm$1.4 85.7$\pm$0.2
Loss Components: Table 3 presents the model’s performance comparison when different loss components are applied. The models utilize pretrained initialization: ResNet with unimodal pretraining and Conformer with SSL. Two key observations can be made:
  1. 1.
    The concurrent exploitation of both $L_{MIPD}$ and $L_{Con}$ is yielding consistent improvement. Exploiting them separately leads to smaller improvement for $L_{Con}$ and even to a decline for $L_{MIPD}$. suggests that alignment in the latent space between the modalities is necessary for the permutations to be effective.
  2. 2.
    The $L_{CEB}$ term, which penalizes task-irrelevant information, improves the Conformer model’s performance, likely due to its pretraining on large, unlabelled datasets that introduce such irrelevant information. In contrast, for the ResNet model, where pretraining already focuses on task-related information, the $L_{CEB}$ term does not provide additional benefits.
Table 3: Ablation Study: Regularization Components – Accuracy (%) of ResNet and Conformer models across datasets with different combinations of $\operatorname{MCR}$ components: $L_{MIPD}$, $L_{Con}$, and $L_{CEB}$. Results indicate that combining $L_{MIPD}$ and $L_{Con}$ is crucial for improvement, while $L_{CEB}$ does not benefit all models.
MCR Components ResNet Conformer
$L_{MIPD}$ $L_{Con}$ $L_{CEB}$ CREMA-D AVE UCF CREMA-D AVE
73.4$\pm$2.5 71.1$\pm$1.4 50.0$\pm$2.0 84.1$\pm$0.6 87.9$\pm$1.1
74.1$\pm$2.9 72.1$\pm$0.5 49.4$\pm$1.9 83.9$\pm$1.8 87.8$\pm$1.7
73.4$\pm$2.1 72.6$\pm$0.6 54.8$\pm$1.2 84.5$\pm$0.3 88.7$\pm$1.4
76.2$\pm$1.7 73.3$\pm$0.5 55.1$\pm$0.6 84.5$\pm$0.3 88.7$\pm$1.4
75.6$\pm$1.9 72.1$\pm$0.9 54.7$\pm$1.1 85.7$\pm$0.2 88.8$\pm$1.0
Figure 5: Error comparison matrices across pairs of datasets and models, comparing unimodal predictions with multimodal models trained using various methods, including Ensemble, Joint Training, AGM, MLB, and MCR. Each column of the confusion matrix represents cases where both unimodal predictions are incorrect, where only one is correct, and where both are correct. The results highlight that MCR consistently performs well in cases where at least one unimodal prediction is correct, with ResNet-based models surpassing even the Ensemble in this category. Additionally, MLB demonstrates stronger performance in discovering emergent information, highlighting a current limitation of MCR in this aspect.
4.4 Analysis of Multimodal Error
In this section, we investigate the specific data points that MCR prioritizes during training. The framework is designed to balance and enhance the independent contributions of each modality through the CMI terms, while strengthening their shared contributions via the contrastive loss. However, no specific term explicitly addresses the synergetic, emergent information that arises from the interplay between modalities. The underlying expectation is that encouraging modalities to improve concurrently would naturally lead to such synergetic behavior. Here, we conduct a post-hoc error analysis to examine the results and better understand the benefits that MCR provides.
We analyze a matrix comparing the unimodal wrong and correct predictions with those from multimodal models trained using various methods, including Ensemble, Joint Training, AGM, MLB, and MCR. As shown in Figure 5, a consistent pattern emerges across different dataset/model pairs. MCR outperforms previous methods in cases where at least one modality independently makes the correct prediction, with the ResNet-based models even surpassing the Ensemble in this category. However, we observe that while MLB and AGM struggle to retain unimodal information effectively, they consistently perform better at discovering new emergent information. MCR improves upon this in certain instances by approaching the best-performing baselines but it generally shows a decline in that category. This analysis is encouraging for two reasons: it highlights a significant percentage of unimodal information that is not adequately leveraged by the multimodal model, and it indicates potential for further improving the discovery of emergent information, both of which are crucial for achieving performance gains that justify the effort required to collect, process, and compute the additional multimodal data.
5 Discussion
This paper examines the challenge of modality competition in multimodal learning, where certain modalities dominate the training process, resulting in suboptimal performance. We introduce the Multimodal Competition Regularizer (MCR), a novel approach inspired by information theory, which frames multimodal learning as a game where each modality competes to maximize its contribution to the final output. MCR efficiently computes lower and upper bounds to optimize both unique and shared task-relevant information for each modality. Our extensive experiments show that MCR consistently outperforms existing methods and simple baselines on both synthetic and real-world datasets, providing a more balanced and effective multimodal learning framework. MCR paves the way for fulfilling the long-standing promise of multimodal fusion methods to achieve performance that surpasses the combined results of unimodal training.
Looking ahead, further advancements could involve incorporating additional terms to explicitly encourage synergetic interactions between modalities, alongside an exploration of game-theoretic modeling to enable more flexible and adaptive strategies. Future work may focus on refining these strategies, enabling individualized and adaptive decisions for each modality, and tightening mutual information bounds to unlock even greater performance improvements.
Appendix A Appendix / supplemental material
A.1 Evidence of limitations of supervised multimodal training
Identifying the instances where supervised multimodal training collapses is often easier than resolving the issue itself. Nevertheless, it is crucial to understand both these limitations and why simple solutions might not suffice. To explore this, we utilize a ResNet-18 backbone on the CREMA-D dataset, employing audio and video as the two modalities. These modalities are concatenated just before the final linear layer. We measure multimodal performance at the end of each epoch and, simultaneously, perform linear probing on each modality to evaluate their individual contributions throughout training.
In Figure 6, it is clear that the performance of the multimodal model aligns closely with that of the audio modality alone, suggesting that the model heavily relies on audio while neglecting the video modality. This lack of exploration results in the video modality remaining at chance-level accuracy throughout training. As a result, the model fails to leverage the any information available in the video modality and performs significantly worse than an ensemble of the unimodally trained models.
A.2 Datasets
@@
CREMA-D [4]: is an emotion recognition dataset with audio and video modalities.It features a diverse group of 91 actors, covering a wide range of ages, ethnicities, and genders. To ensure consistency, each actor is positioned at an equal distance from the camera, expressing six distinct emotions: Happy, Sad, Anger, Fear, Disgust, and Neutral. In alignment with methodologies from prior studies [40; 28; 6], video frames are sampled at 1 fps, selecting 3 consecutive frames, while audio segments are sampled at 22 kHz, capturing 3 seconds that correspond with the chosen video frames. Audio analysis utilizes a window size of 512 and a step size of 353 samples for Short-Time Fourier Transform (STFT), creating log-Mel spectrograms. For advanced models, our methodology aligns with Goncalves et al. [13], incorporating audio signals sampled at 16 kHz and utilizing pre-calculated facial features. Unlike previous approaches [40; 28; 6], our dataset division follows Goncalves et al. [13], excluding actor overlap between training, validation, and test sets. We report standard deviation (std) across folds for consistency.
@@
Something-Something (V2)[15]: presents 220,847 video clips where individuals execute 174 distinct, object-agnostic hand actions, spanning a wide range of simple hand movements without reliance on specific objects. This extensive collection markedly exceeds the data volume of prior datasets. In alignment with Radevski et al. [43], we integrate Optical Flow (OF) as the additional modality for capturing dynamic motions. We sample videos at 1 fps, retaining 16 frames per sample. Our data preparation adheres to the pipeline outlined in Radevski et al. [43]. Results are reported based on a single run.
@@
CMU-MOSI[61] and -MOSEI[63]: datasets serve as benchmarks for multimodal sentiment and emotion analysis. MOSI comprises 2,199 video clips with video, audio and text modalities. MOSEI expands on this with around 10x more YouTube movie review clips from 1000 different speakers. Both datasets are annotated with sentiment scores (-3 to 3) and following previous works [50] we employ metrics such as 7-class accuracy, binary accuracy, F1 score, mean absolute error, and correlation with human annotations for model evaluation. We use for both datasets the aligned versions following [30].
A.3 Models: Backbone Unimodal Encoders
@@
In line with previous research [28; 40; 6; 59], our initial experiments adopt ResNet-18 [18] as the unimodal encoder for handling both video and audio modalities in the CREMA-D, AVE, and UCF datasets. These models are randomly initialized and incorporate adaptive pooling to accommodate diverse input dimensions. For the CMU-MOSEI dataset, we exploit a 5-layer Transformer [51] similar to previous works [31; 30].
@@
We extend our investigation to include a larger, pre-trained set of unimodal encoders that are optimally suited for each specific modality. This selection process is designed to rigorously assess whether state-of-the-art models exhibit susceptibility to the same phenomena under investigation. We exploit these encoders on the CREMA-D dataset. Following [14] on CREMA-D, we deploy the first 12 layers of the Wav2Vec2 [3; 56] model with self-supervised pretrained weights for speech recognition, allowing the Wav2Vec2 model to be finetuned. For the video modality we extract the facing bounding boxes with the multi-task cascaded convolutional neural network (MTCNN) face detection algorithm [64] and afterwards the facial features of every available frame exploiting EfficientNet-B2 [48] as a frozen feature descriptor. The extracted audio features and pre-calculated facial features are further refined using a 5-layer Conformer [16], initialized from scratch. Although the model includes additional components, we refer to it as ’Conformer’ for simplicity.
@@
For the AVE dataset, we use a similar architecture to CREMA-D, where each branch utilizes an advanced, pretrained model aligned with AVE data, followed by a 5-layer Conformer. For the video branch, we use the ViViT model [2] pretrained on Kinetics [23], and for the audio branch, the HuBERT [19] model pretrained on Audioset [12]. Ensure the audio pretraining included non-speech data was important for the pretraining to be beneficial. We also refer to this model as ”Conformer,” although it incorporates different large pretrained models in this instance.
@@
In the case of the Something-Something dataset, our methodology builds upon the insights presented by Radevski et al. [43], which highlight the importance of modality-specific processing in multimodal tasks. For both video and optical flow data, we adopt the Swin Transformer [33] as backbone to each modality. This state-of-the-art architecture excels in capturing hierarchical and spatiotemporal features through its shifted window attention mechanism.
A.4 Experimental Details
@@
In this section, we outline the necessary details to reproduce our experiments. Across all datasets, we follow a consistent procedure: we first determine an appropriate learning rate (lr) for the unimodal models by testing several candidates until finding one that works across both modalities. While this step could be avoided by exploiting parameter specific learning rates, we expected stability implications which we aimed to avoid. For all the experiments of the same dataset/model pair we use the same hyperparameters, except when fine-tuning pretrained encoders, where we apply a learning rate scaled by one magnitude lower.
@@
All models are optimized using Adam [25] with a cosine learning rate scheduler and a steady warm-up phase, except for the Something-Something dataset, where we use Adaw [34]. Early stopping is applied for all models, with maximum epochs set to 100 for ResNet and Transformer models, 50 for Conformer models, and 30 epochs in total for Swin Transformers without early stopping. Batch sizes are adjusted based on computational resources, with ResNets and Transformers both using a batch size of 32, Conformers using 8, and Swin-TF using 16. These settings ensure balanced performance and efficient training across all experiments. We use different learning rates ($lr$) and weight decay ($wd$) values across experiments, tailored to each dataset and model. For ResNet models, we use $lr=1\text{e}{-3}$ and $wd=1\text{e}{-4}$ for CREMA-D, while both AVE and UCF use $lr=1\text{e}{-4}$ and $wd=1\text{e}{-4}$. For Transformer models, including MOSI and MOSEI on two and three modalities, the hyperparameters are consistently $lr=1\text{e}{-4}$ and $wd=1\text{e}{-4}$. Similarly, for Conformer models, we set $lr=5\text{e}{-5}$ and $wd=5\text{e}{-6}$ for CREMA-D, while AVE uses $lr=1\text{e}{-4}$ and $wd=1\text{e}{-4}$. Finally, for Swin-TF models trained on the Something-Something dataset, we configure $lr=1\text{e}{-4}$ and $wd=0.02$.
@@
Each of the previous methods includes its own set of hyperparameters, typically just one, with some exceptions such as MSLR, which requires additional parameters. For each dataset/model combination, we conduct a brief hyperparameter search, ensuring an equitable number of trials across methods. Due to the extensive list of hyperparameters, we will provide detailed configurations for each experiment in our GitHub repository. The repository link will be included here following the double-blind review process.
A.5 MCR on M modalities

### v3

@@ 4.2 Real-World Datasets
    $\operatorname{MCR}$ is the only method that consistently surpasses all baselines across datasets, model architectures, number of modalities, and task types (classification and regression), with an exception on AVE with Conformer model. This consistent advantage highlights $\operatorname{MCR}$’s ability to balance modality contributions during training under diverse settings, positioning it as a strong and generalizable approach for multimodal learning.
4.3 Analysis of Multimodal Error
Methodology: To understand how $\operatorname{MCR}$ improves over existing methods, we perform a post-hoc error analysis by categorizing each sample based on the correctness of the unimodal predictions. Specifically, we consider four groups: (1) both unimodal models are correct, (2) only the first is correct, (3) only the second is correct, and (4) both are incorrect. This breakdown allows us to compare how different multimodal methods behave across these categories and identify whether gains arise from selective reliance on unimodal cues or from synergistic integration.
Results: The error analysis of Figure 4 reveals key strengths and limitations of the proposed method. $\operatorname{MCR}$ consistently outperforms other methods in routing information favouring both modalities in the cases that only one of the modalities is correct maintaining competitive performance on both of them and in the case that all modalities correctly predict the label. This demonstrates $\operatorname{MCR}$’s ability to effectively route information to the appropriate modality, in line with its design choice to model and control training via this modality-independent, task-relevant information through the mutual information terms.
However, $\operatorname{MCR}$ does not exhibit significant gains in capturing synergetic information in the datapoints that all unimodal models fail, underperforming relative to AGM and MLB. This suggests $\operatorname{MCR}$ excels in routing decisions but may be less effective at leveraging synergies across modalities when all individual models falter. This trend holds across other datasets (see Appendix A.11), with the exception of MOSI, where $\operatorname{MCR}$ improves synergy. Our initial assumption that concurrent modality training would foster synergy thus did not hold in practice.
Figure 4: Error comparison on the CREMA-D dataset across unimodal and multimodal models ($\operatorname{MCR}$, Ensemble, Joint Training, AGM, MLB). Each matrix summarizes model performance based on unimodal prediction correctness. $\operatorname{MCR}$ performs best when at least one unimodal branch is correct (brown box), effectively preserving modality-specific signals. However, AGM and MLB outperform $\operatorname{MCR}$ when both unimodal predictions fail, in the "Both Wrong" (purple box), indicating stronger synergy in those edge cases. Trends across other datasets are shown in Appendix A.11, with MOSI being a notable exception where $\operatorname{MCR}$ also excels in synergy.
5 Discussion
This paper examines the challenge of modality competition in multimodal learning, where certain modalities dominate the training process, resulting in suboptimal performance. We introduce the Multimodal Competition Regularizer ($\operatorname{MCR}$), a novel approach inspired by information theory, which frames multimodal learning as a game where each modality competes to maximize its contribution to the final output. $\operatorname{MCR}$ efficiently computes lower and upper bounds to optimize both unique and shared task-relevant information for each modality. Our extensive experiments show that $\operatorname{MCR}$ consistently outperforms existing methods and simple baselines on both synthetic and real-world datasets, providing a more balanced and effective multimodal learning framework. $\operatorname{MCR}$ paves the way for fulfilling the long-standing promise of multimodal fusion methods to achieve performance that surpasses the combined results of unimodal training.
We explored different game strategies and observed that directly encouraging competition between modalities in the overall objective function positively impacts performance, as detailed in Appendix A.8. Future work could investigate more refined strategies to enable individualized and adaptive decisions for each modality to unlock greater performance gains.
Lastly, we conduct a post-hoc error analysis found both in Section 4.3 and Appendix A.11, examining overlaps between the errors of multimodal models and their unimodal counterparts. The results show that $\operatorname{MCR}$ excels at routing decisions to the correct unimodal information but does not promote synergetic behavior accordingly, compared to previous methods. Our initial assumption that the simultaneous progress of unimodal encoders during training would naturally enhance synergy was not supported in practice, highlighting the need for future work to promote this behavior explicitly. Finally, this analysis highlights the potential for performance improvements through enhanced multimodal training, motivating further exploration in this area.
Appendix A Supplementary Material
A.1 Evidence of limitations of supervised multimodal training
Identifying the instances where supervised multimodal training collapses is often easier than resolving the issue itself. Nevertheless, it is crucial to understand both these limitations and why simple solutions might not suffice. To explore this, we utilize a ResNet-18 [50] backbone on the CREMA-D dataset, employing audio and video as the two modalities. These modalities are concatenated just before the final linear layer. We measure multimodal performance at the end of each epoch and, simultaneously, perform linear probing on each modality to evaluate their individual contributions throughout training.
In Figure 5, it is clear that the performance of the multimodal model aligns closely with that of the audio modality alone, suggesting that the model heavily relies on audio while neglecting the video modality. This lack of exploration results in the video modality remaining at chance-level accuracy throughout training. As a result, the model fails to leverage any information available in the video modality and performs significantly worse than an ensemble of the unimodally trained models.
A.2 Multimodal Competition Error
Multimodal competition occurs when the network primarily optimizes for one modality, leading to a decline in generalization. One modality dominates the reduction in training error and limits gradient feedback to the other modalities. Huang et al. [14] proved that in late fusion models (i.e. $\theta_{c}=\emptyset$), each modality has a probability of causing multimodal competition. They define the phenomenon in terms of the correlation $\Gamma_{m}=\sum_{cl}\max[\langle X^{cl}_{m},W^{cl}\rangle]^{+}$ of each input modality and the weights, $W$, of the output layer for each corresponding class $cl$. If a modality dominates the competition, it causes the other modality to maintain its initial (i.e. before training) correlation levels. Such correlation levels do not have a specific target, unlike other problems where equality is the goal [51], making the solution less straightforward. We proceed with this analysis on late-fusion models by introducing the generalization error $\epsilon$ resulting from the effects of multimodal competition.
Definition A.1.
Definition of Multimodal Competition Error (MCE): Let $\Gamma_{1}$ and $\Gamma_{2}$ represent the correlations of modalities $Z_{1}$ and $Z_{2}$ with the output. If, after training, $\Gamma_{1}\gg\Gamma_{2}$, and modality $X_{2}$ has predictive power greater than randomness ($R(X_{2})<R(\text{random})$), while making errors independent from those of $X_{1}$ ($\mathbb{E}[e_{1}\cdot e_{2}]=0$, where $e_{1}$ and $e_{2}$ are the errors of $X_{1}$ and $X_{2}$, respectively), then for a trained model $f$ and its optimal solution $f^{*}$, there exists a multimodal competition error $\epsilon$, such that:
@@ A.2 Multimodal Competition Error
where $R(\cdot)=\mathbb{E}[\mathcal{L}_{\text{task}}(Y,\cdot)]$ and $R_{\text{emp}}(\cdot)$ denote the generalization and empirical model risks, and $\mathcal{L}_{\text{task}}$ represents the corresponding task loss, which could vary depending on the objective, such as cross-entropy for classification or mean-squared error for regression. A higher $\epsilon$ indicates a stronger effect of multimodal competition, implying that the dominance of one modality significantly impacts the model’s generalization. These inequalities are empirically observed through improvements in generalization achieved by adjusting the training objectives to address multimodal competition, without altering the model or the data. Numerically estimating $\epsilon$ would require knowledge of the optimal solution $f^{*}$, which is typically unavailable. Lastly, transitioning from the late-fusion models of Huang et al. [14] we need to extend this definition by replacing correlation with $\operatorname{MI}$ to consider also non-linear statistical dependencies.
A.3 Datasets
@@
CREMA-D [32]: is an emotion recognition dataset with audio and video modalities. It features a diverse group of 91 actors, covering a wide range of ages, ethnicities, and genders. To ensure consistency, each actor is positioned at an equal distance from the camera, expressing six distinct emotions: Happy, Sad, Anger, Fear, Disgust, and Neutral. In alignment with methodologies from prior studies [22, 21, 16], video frames are sampled at 1 fps, selecting 3 consecutive frames, while audio segments are sampled at 22 kHz, capturing 3 seconds that correspond with the chosen video frames. Audio analysis utilizes a window size of 512 and a step size of 353 samples for Short-Time Fourier Transform (STFT), creating log-Mel spectrograms. For advanced models, our methodology aligns with Goncalves et al. [52], incorporating audio signals sampled at 16 kHz and utilizing pre-calculated facial features. Unlike previous approaches [22, 21, 16], our dataset division follows Goncalves et al. [52], excluding actor overlap between training, validation, and test sets. We report standard deviation (std) across folds for consistency.
@@
Something-Something (V2) [35]: presents 220,847 video clips where individuals execute 174 distinct, object-agnostic hand actions, spanning a wide range of simple hand movements without reliance on specific objects. This extensive collection markedly exceeds the data volume of prior datasets. In alignment with Radevski et al. [5], we integrate Optical Flow (OF) as the additional modality for capturing dynamic motions. We sample videos at 1 fps, retaining 16 frames per sample. Our data preparation adheres to the pipeline outlined in Radevski et al. [5]. Results are reported based on three random seeds on the same validation set.
@@
CMU-MOSI [33] and -MOSEI [34]: datasets serve as benchmarks for multimodal sentiment and emotion analysis. MOSI comprises 2,199 video clips with video, audio and text modalities. MOSEI expands on this with around 10x more YouTube movie review clips from 1000 different speakers. Both datasets are annotated with sentiment scores (-3 to 3) and following previous works [8] we employ metrics such as 7-class accuracy, binary accuracy, F1 score, mean absolute error, and correlation with human annotations for model evaluation. We use for both datasets the aligned versions following [53]. The reported std is derived from three random seeds on the same test set.
A.4 Models: Backbone Unimodal Encoders
@@
In line with previous research [21, 22, 16, 40], our initial experiments adopt ResNet-18 [50] as the unimodal encoder for handling both video and audio modalities in the CREMA-D, AVE, and UCF datasets. These models are randomly initialized and incorporate adaptive pooling to accommodate diverse input dimensions. For the CMU-MOSEI dataset, we exploit a 5-layer Transformer [54] similar to previous works [41, 53].
@@
We extend our investigation to include a larger, pre-trained set of unimodal encoders that are optimally suited for each specific modality. This selection process is designed to rigorously assess whether state-of-the-art models exhibit susceptibility to the same phenomena under investigation. We exploit these encoders on the CREMA-D dataset. Following [49] on CREMA-D, we deploy the first 12 layers of the Wav2Vec2 [55, 56] model with self-supervised pretrained weights for speech recognition, allowing the Wav2Vec2 model to be finetuned. For the video modality, we extract the facing bounding boxes with the multi-task cascaded convolutional neural network (MTCNN) face detection algorithm [57] and afterward the facial features of every available frame exploiting EfficientNet-B2 [58] as a frozen feature descriptor. The extracted audio features and pre-calculated facial features are further refined using a 5-layer Conformer [59], initialized from scratch. The total model size is 183M parameters. Although the model includes additional components, we refer to it as "Conformer" for simplicity.
@@
For the AVE dataset, we use a similar architecture to CREMA-D, where each branch utilizes an advanced, pretrained model aligned with AVE data, followed by a 5-layer Conformer. For the video branch, we use the ViViT model [60] pretrained on Kinetics [61], and for the audio branch, the HuBERT [62] model pretrained on Audioset [63] resulting in a model with 232M parameters. Ensuring the audio pretraining included non-speech data was important for the pretraining to be beneficial. We also refer to this model as "Conformer," although it incorporates different large pretrained models in this instance.
@@
In the case of the Something-Something dataset, our methodology builds upon the insights presented by Radevski et al. [5], which highlight the importance of modality-specific processing in multimodal tasks. For both video and optical flow data, we adopt the Swin Transformer [48] as the backbone encoder to each modality. This state-of-the-art architecture excels in capturing hierarchical and spatiotemporal features through its shifted window attention mechanism.
A.5 Experimental Details
@@
In this section, we outline the necessary details to reproduce our experiments. Across all datasets, we follow a consistent procedure: we first determine an appropriate learning rate (lr) for the unimodal models by testing several candidates until finding one that works across both modalities. While this step could be avoided by exploiting parameter-specific learning rates, we expected stability implications which we aimed to avoid. For all the experiments of the same dataset/model pair, we use the same hyperparameters, except when fine-tuning pretrained encoders, where we apply a learning rate scaled one magnitude lower.
@@
All models are optimized using Adam [64] with a cosine learning rate scheduler and a steady warm-up phase, except for the Something-Something dataset, where we use Adaw [65]. Early stopping is applied for all models, with maximum epochs set to 100 for ResNet and Transformer models, 50 for Conformer models, and 30 epochs in total for Swin Transformers without early stopping. Batch sizes are adjusted based on computational resources, with ResNets and Transformers both using a batch size of 32, Conformers using 8, and Swin-TF using 16. These settings ensure balanced performance and efficient training across all experiments. In several instances, initializing the encoders with pre-trained weights from unimodal training proved beneficial. This was particularly effective for datasets without precomputed features and models without access to larger-scale SSL pretraining. We use different learning rates ($lr$) and weight decay ($wd$) values across experiments, tailored to each dataset and model. For ResNet models, we use $lr=1\text{e}{-3}$ and $wd=1\text{e}{-4}$ for CREMA-D, while both AVE and UCF use $lr=1\text{e}{-4}$ and $wd=1\text{e}{-4}$. For Transformer models, including MOSI and MOSEI on two and three modalities, the hyperparameters are consistent $lr=1\text{e}{-4}$ and $wd=1\text{e}{-4}$. Similarly, for Conformer models, we set $lr=5\text{e}{-5}$ and $wd=5\text{e}{-6}$ for CREMA-D, while AVE uses $lr=1\text{e}{-4}$ and $wd=1\text{e}{-4}$. Finally, for Swin-TF models trained on the Something-Something dataset, we configure $lr=1\text{e}{-4}$ and $wd=0.02$.
@@
Each of the previous methods includes its own set of hyperparameters, typically just one, with some exceptions such as MSLR or D&R, which requires additional parameters. For each dataset/model combination, we conduct a brief hyperparameter search, ensuring an equitable number of trials across methods. Due to the extensive list of hyperparameters, we will provide detailed configurations for each experiment in our GitHub repository. The repository link will be included here following the double-blind review process.
Lastly, all of our experiments run on single GPU with different nodes being used for different experiments. For the largest ones we utilized H100 with 80Gb vram to run the experiments of Sth-Sth which required the longest of all up to 48 hours per run. For the rest, we would have from some minutes on the smallest experiment up to 4-5 hours for the datasets CREMA-D, AVE and UCF depending on the GPU and the available RAM.
A.6 Bounding task-irrelevant information via CEB
Our objective is to maximize the information shared between modalities that is irrelevant to the supervised task. Directly estimating the conditional mutual information $I(X_{1};X_{2}\mid Y)$ is challenging in high-dimensional settings, so we relax the objective by penalizing both the individual and shared irrelevant information, leading to the decomposition:
$I(X_{1};X_{2}\mid Y)=H(X_{1}\mid Y)+H(X_{2}\mid Y)-H(X_{1},X_{2}\mid Y)$ (15)
$\Rightarrow-I(X_{1};X_{2}\mid Y)+H(X_{1}\mid Y)+H(X_{2}\mid Y)=H(X_{1},X_{2}\mid Y)$ (16)
We lower-bound the conditional joint entropy using CEB [47] as follows,
$H(X_{1},X_{2}\mid Y)$ $=-\mathbb{E}_{p(x_{1},x_{2},y)}\left[\log p(x_{1},x_{2}\mid y)\right]$ (17)
$=-\mathbb{E}_{p(x_{1},x_{2},y)}\left[\log g(x_{1},x_{2}\mid y)\right]-\mathrm{KL}\big(p(x_{1},x_{2}\mid y)\,\|\,g(x_{1},x_{2}\mid y)\big)$ (18)
$\leq-\mathbb{E}_{p(x_{1},x_{2},y)}\left[\log g(x_{1},x_{2}\mid y)\right]$ (19)
Assuming a conditional Gaussian model $g(x_{1},x_{2}\mid y)=\mathcal{N}((x_{1},x_{2});\mu(y),\sigma^{2}I)$, we define $\mu(y)$ as a deterministic function $h:Y;\theta_{h}\rightarrow Z=(Z_{1},Z_{2})$ that predicts a target joint representation from $Y$. Then, the conditional entropy term is upper-bounded as:
$H(X_{1},X_{2}\mid Y)\leq-\mathbb{E}_{p(x_{1},x_{2},y)}\left[\log g(x_{1},x_{2}\mid y)\right]\propto\mathbb{E}_{p(x_{1},x_{2},y)}\left\|[f(x_{1}),f(x_{2})]-h(y;\theta_{h})\right\|^{2}$ (20)
Thus, we approximate the entropy term with an MSE loss, encouraging $(x_{1},x_{2})$ to deviate from any deterministic function when it is not informative for the task. Maximizing $H(X_{1}\mid Y)$ and $H(X_{2}\mid Y)$ incentivizes each modality to retain information that is task-independent, aligning with the goal of minimizing $I(X_{1};X_{2}\mid Y)$ and suppressing task-irrelevant alignment between modalities.
A.7 MCR on M modalities